OpenStudy (idealist10):

Find solutions y1 and y2 of the equation y"=0 that satisfy the initial conditions y1(xo)=1, y'1(xo)=0 and y2(xo)=0, y'2(xo)=1.

OpenStudy (idealist10):

@ganeshie8 @ParthKohli @iambatman @zepdrix

OpenStudy (idealist10):

@nincompoop @pooja195 @Directrix @Nnesha

OpenStudy (idealist10):

Wait, you said to integrate y"=0, right?

zepdrix (zepdrix):

Ya that would be a good first step :) To integrate it twice. I'm still thinking about this one though XD

zepdrix (zepdrix):

It looks like you asked this question 6 months ago lol I'm looking at it currently, seeing what was discussed.

OpenStudy (idealist10):

\[y'=c\]

OpenStudy (idealist10):

\[y=c _{1}+c _{2}x\]

OpenStudy (idealist10):

Now what? What do I do?

zepdrix (zepdrix):

K cool. I guess we could say this: \(\large\rm y=y_1+y_2\) So we can write our equation like this: \(\large\rm y_1+y_2=c_1+c_2 x\) Differentiating gives us: \(\large\rm y_1'+y_2'=c_2\)

zepdrix (zepdrix):

And then use the initial data to figure out some stuff.

zepdrix (zepdrix):

Given that, \(\large\rm \color{royalblue}{y_1(x_o)=1}\) and \(\large\rm \color{orangered}{y_2(x_o)=0}\) We have:\[\large\rm \color{royalblue}{y_1(x_o)}+\color{orangered}{y_2(x_o)}=\color{royalblue}{1}+\color{orangered}{0}=c_1+c_2 x_o\]

zepdrix (zepdrix):

So that gives us an equation for our constants,\[\large\rm 1=c_1+c_2 x_o\]

zepdrix (zepdrix):

Also given, \(\large\rm \color{royalblue}{y_1'(x_o)=0}\) and \(\large\rm \color{orangered}{y_2'(x_o)=1}\) We have:\[\large\rm \color{royalblue}{y_1'(x_o)}+\color{orangered}{y_2'(x_o)}=\color{royalblue}{0}+\color{orangered}{1}=c_2\]Which gives us\[\large\rm c_2=1\]

zepdrix (zepdrix):

Plug this information back into the other constant equation,\[\large\rm 1=c_1+c_2 x_o\qquad\to\qquad 1=c_1+1x_o\]Solving for c_1 gives us,\[\large\rm c_1=1-x_o\]And recall that we also have:\[\large\rm c_2=1\]

zepdrix (zepdrix):

If we go all the way back to our equation:\[\large\rm y_1+y_2=c_1+c_2 x\]we have:\[\large\rm y_1+y_2=(1-x_o)+1 x\]

zepdrix (zepdrix):

And it looks like the last time you asked this question, you had said that the solutions are supposed to be: \(\large\rm y_1=1,\qquad y_2=x-x_0\) Is that still the case? Do you have access to the answer? Seems like the solutions should be: \(\large\rm y_1=1-x_o,\qquad y_2=x\) Hmm..

OpenStudy (idealist10):

Yes, well, the book has the answers y1=1, y2=x-x0. But are you really sure that your answers are correct? The book answers are wrong?

zepdrix (zepdrix):

No, I'm not confident enough to say that the book is wrong, I'm a little rusty on my ODE's. :) I could be wrong, seems more likely that the book answer is incorrect though. That's what Sith had suggested last time as well it seems.

OpenStudy (idealist10):

Dude, I finally found the way to do this problem and the book's answers are right! Thank you for the help though. :)

zepdrix (zepdrix):

Oh interesting :OO Now I'm kind of curious lol